One number exceeds another by 24. The sum of the numbers is 58. What are the numbers?
Answer by checkley77(12844)

y=x+24
x+y=58
x+x=24=58
2x=58-24
2x=34
x=34/2
x=17 ans.
y=17+24=41 ans.
Proof:
17+41=58
58=58